# Do We Really Need Permutations? Impact of Width Expansion on Linear Mode Connectivity

### v1

@@ 5.3 Empirical Verification of Reciprocal Orthogonality
From the above, we confirm that reciprocal orthogonality holds in wide models. This is also attributable to the low-rank structure of the weights. For example, if a trained model’s weight matrices have high rank, then since one model’s weight $\bm{W}_{\ell}^{(a)}$ at some layer $\ell$ has high rank, $\|\bm{W}_{\ell}^{(a)}\bm{z}_{\ell-1}\|/\|\bm{z}_{\ell-1}\|$ becomes large regardless of the direction of the input $\bm{z}_{\ell-1}$. This implies that reciprocal orthogonality does not hold. Ito et al. (2025a) clarify that the relative rank of weights with respect to width becomes smaller for wider models, explaining why reciprocal orthogonality is more likely to hold when the model is wider.
@@ 5.4 Effect of Weight Decay on LMC
Prior works (Timor et al., 2023; Galanti et al., 2022; Ito et al., 2025a) empirically observed that widening models and stronger weight decay encourage low-rank weight matrices. Conversely, with weak weight decay, weight matrices tend to have higher rank, making LEWC less likely. Figure 9 shows that merged models trained with weak weight decay indeed exhibit large barriers at $\lambda=0.5$. In Section E.5, we also demonstrate that weak additivity and reciprocal orthogonality fail in this setting. These results underscore that low-rank structure is crucial for LMC through LEWC.
@@ 6 Conclusion
While LMC is broadly relevant, our analysis focused on standard image classification. Because LEWC typically requires larger width multiplier than permutation-based merging, we limited experiments to relatively simple architectures and datasets. An important direction for future work is to test whether these phenomena persist in large-scale settings and other modalities.
@@
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@
The experimental settings for reproducing the results are described in Appendix B. All proofs of the theorems are given in Appendix D.
@@ Appendix A Additional Related Work
A series of works (Garipov et al., 2018; Draxler et al., 2018; Freeman & Bruna, 2017) have shown that independently trained neural networks can often be connected by low-loss nonlinear paths. Nagarajan & Kolter (2019) provided an early observation that, on MNIST, solutions obtained from stochastic gradient descent (SGD) with the same initialization can be connected by linear paths with little increase in loss. Later, Frankle et al. (2020) systematically demonstrated that such linear connections are not universal: whether they appear depends strongly on the dataset and architecture. They also observed that when two models are branched from a shared partially trained model, the resulting solutions are almost always connected by a linear path. Moreover, they investigated the link between linear mode connectivity (LMC) and the lottery ticket hypothesis (Frankle & Carbin, 2019).
More recent work has explored symmetry and alignment as explanations for LMC. Entezari et al. (2022) conjectured that, once permutation symmetries of hidden units are taken into account, LMC should hold with high probability. Building on this, Ainsworth et al. (2023) introduced a weight-matching approach based on bipartite graph alignment, while Guerrero-Peña et al. (2023) applied Sinkhorn’s algorithm as a relaxation to solve the alignment problem more directly. Although a number of papers (Venturi et al., 2019; Nguyen et al., 2019; Nguyen, 2019; Kuditipudi et al., 2019) have studied nonlinear mode connectivity, theoretical understanding of LMC remains limited. Ferbach et al. (2024) established width-dependent conditions under which LMC can be guaranteed, assuming independence of weight vectors. Zhou et al. (2023) proposed the notion of layerwise linear feature connectivity (LLFC) and proved that LLFC implies LMC. More recently, Ito et al. (2025a) highlighted the role of dominant singular vectors in parameter space, showing that they are critical to satisfying LMC, especially when alignment via weight matching is applied.
In addition, several theoretical works have analyzed the effect of network width on nonlinear connectivity. For instance, Nguyen (2019) proved that in over-parameterized neural networks with piecewise-linear activations, every sublevel set of the loss is connected and unbounded. Shevchenko & Mondelli (2020) further showed that in wide multilayer perceptrons, SGD solutions can be linked through piecewise-linear paths along which the loss barrier vanishes as width increases. However, to the best of our knowledge, there are no theoretical results directly establishing that larger width makes LMC more likely. This motivates our empirical investigation of that question.
@@
Model merging has been studied in close connection with LMC, as well as in the context of federated learning. The idea of federated learning was introduced by McMahan et al. (2017) and Konečný et al. (2016), where models are trained locally on partitioned datasets. Wang et al. (2020) explored permuting local model components prior to aggregation, while Singh & Jaggi (2020) proposed an approach to merge models by optimal transport, conceptually related to the weight-matching method of Ainsworth et al. (2023). Although Singh et al. (2024)’s method was primarily designed for fusion and empirically underperforms the latter, it can still be interpreted as an LMC-based technique because it enforces alignment within the same architecture.
While our work focuses on merging models trained from different random seeds, there is a parallel line of research on merging models obtained by fine-tuning from the same fundamental model. A representative example is Model Soups (Wortsman et al., 2022), which showed that averaging weights across fine-tuned models trained under different hyperparameters can improve test accuracy without additional inference cost, contrasting with standard ensembling. Matena & Raffel (2022) extended this idea by weighting models according to the Fisher information matrix, leading to more effective combinations. Yadav et al. (2023) introduced TIES-Merging, which prunes parameters with small updates, resolves conflicting signs, and averages only aligned components. These approaches are primarily designed for merging fine-tuned models of the same pretrained model, not for models independently trained from scratch with different initializations. By contrast, our study addresses this more challenging setting and seeks to shed light on the feasibility of model merging under such conditions.
@@ Appendix B Experimental Settings
This section details the setup used to train neural networks and obtain solutions from stochastic gradient descent (SGD). The experiments were conducted on four benchmark datasets: MNIST (Lecun et al., 1998), Fashion-MNIST (FMNIST) (Xiao et al., 2017), CIFAR-10 (Krizhevsky & Hinton, 2009), and CIFAR-100. For weight matching (WM), we adopted the implementation available in the public repository of Ainsworth et al. (2023) [^2] https://github.com/samuela/git-re-basin.
All training and evaluation procedures were performed on a Linux workstation equipped with two AMD EPYC 7543 32-core processors, eight NVIDIA A30 GPUs, and 512 GB of memory. We used PyTorch 2.5.1333https://pytorch.org/, PyTorch Lightning 2.4.0444https://lightning.ai/docs/pytorch/stable/, and torchvision 0.20.1555https://pytorch.org/vision/stable/index.html as the software framework.
@@ B.1 Model Training
In line with the setup of Ainsworth et al. (2023), we trained a fully connected multilayer perceptron (MLP) with three hidden layers, each containing 512 units. ReLU activations were applied to the hidden layers. For both MNIST and FMNIST, training used the Adam optimizer with a fixed learning rate of $1\times 10^{-3}$ and a weight decay of $3\times 10^{-3}$. The batch size was 512, and models were trained for up to 100 epochs. No learning rate scheduling was applied.
@@
For these experiments, we employed the source code released by Ito et al. (2025b) [^6] https://github.com/e5-a/STE-MM. The VGG-11 and ResNet-20 architectures followed the implementations described by Ainsworth et al. (2023). During model merging, BatchNorm statistics were re-calibrated using the training data, following the procedure of Jordan et al. (2023). Models were optimized with SGD using a learning rate of $0.4$ and a weight decay of $3\times 10^{-3}$. The batch size was set to 500, and training was carried out for a maximum of 100 epochs. Data augmentation included random $32\times 32$ crops and random horizontal flips.
@@ Appendix C Layerwise Linear Feature Connectivity
Zhou et al. (2023) pointed out that the existence of LMC in settings such as permutation and spawning can be attributed to a property called layerwise linear feature connectivity (LLFC). Additionally, Zhou et al. (2023) also showed that two conditions are sufficient for LLFC to hold: weak additivity for ReLU activations (Definition 5.1) and the commutativity property defined as follows:
@@
Appendix D Proof
D.1 Main Theorem
Theorem D.1.

### v2

@@ 5.3 Empirical Verification of Reciprocal Orthogonality
From the above, we confirm that reciprocal orthogonality holds in wide models. This is also attributable to the low-rank structure of the weights. For example, if a trained model’s weight matrices have high rank, then since one model’s weight $\bm{W}_{\ell}^{(a)}$ at some layer $\ell$ has high rank, $\|\bm{W}_{\ell}^{(a)}\bm{z}_{\ell-1}\|/\|\bm{z}_{\ell-1}\|$ becomes large regardless of the direction of the input $\bm{z}_{\ell-1}$. This implies that reciprocal orthogonality does not hold. Ito_ICLR_2025 clarify that the relative rank of weights with respect to width becomes smaller for wider models, explaining why reciprocal orthogonality is more likely to hold when the model is wider.
@@ 5.4 Effect of Weight Decay on LMC
Prior works (Timor_CoLT_2023; Tomer_arxiv_2023; Ito_ICLR_2025) empirically observed that widening models and applying stronger weight decay encourage low-rank weight matrices. Conversely, when weight decay is weak, weight matrices tend to have higher rank, making LEWC less likely. Figure 9 shows that merged models trained with weak weight decay indeed exhibit large barriers. To understand this phenomenon in more detail, we examine whether LEWC and its sufficient conditions still hold in this setting. The empirical results are summarized in Figure 10. The impact of weaker weight decay is especially pronounced for VGG-11 and ResNet-20. First, Figure 10(a) indicates that LEWC is no longer satisfied in deeper layers when $\lambda=1/2$. Moreover, Figures 10(b) and 10(c) show that neither weak additivity of ReLU activations nor reciprocal orthogonality holds. Overall, these results show that weakening weight decay breaks LEWC and its two sufficient conditions, highlighting the role of low-rank structure in enabling LMC.
(b) Average cosine similarity between $\sigma((\tilde{\bm{z}}_{\ell}^{(a)}+\tilde{\bm{z}}_{\ell}^{(b)})/2)$ and $(\sigma(\tilde{\bm{z}}_{\ell}^{(a)})+\sigma(\tilde{\bm{z}}_{\ell}^{(b)}))/2$.
Figure 10: Experimental results for trained models with weaker weight decay of $10^{-4}$. From top to bottom, the figures show evaluations of LEWC, weak additivity, and reciprocal orthogonality.
@@ 6 Conclusion
Our experiments focused on standard image classification with relatively simple architectures, since LEWC typically requires larger width multipliers than permutation-based merging. An important direction for future work is to test whether these phenomena extend to larger-scale models and other modalities.
Our work is analytical rather than a proposal of a merging or federated learning method. By identifying weak additivity and reciprocal orthogonality as the drivers of permutation-free LMC, we point to practical directions for developing new model-merging and federated learning methods, such as training or permutation-search procedures that promote these properties, and simple rescaling to offset any LEWC-induced logit norm decay.
@@
This paper presents work whose goal is to advance the field of machine learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
@@
The experimental settings for reproducing the results are described in Appendix B. All proofs of the theorems are given in Appendix E.
@@ Appendix A Additional Related Work
A series of works (Garipov_NIPS_2018; Draxler_ICML_2018; Freeman_ICLR_2017) have shown that independently trained neural networks can often be connected by low-loss nonlinear paths. Nagarajan_NIPS_2019 provided an early observation that, on MNIST, solutions obtained from stochastic gradient descent (SGD) with the same initialization can be connected by linear paths with little increase in loss. Later, frankle_ICML_2020 systematically demonstrated that such linear connections are not universal: whether they appear depends strongly on the dataset and architecture. They also observed that when two models are branched from a shared partially trained model, the resulting solutions are almost always connected by a linear path. Moreover, they investigated the link between linear mode connectivity (LMC) and the lottery ticket hypothesis (Frankle_ICLR_2019).
More recent work has explored symmetry and alignment as explanations for LMC. Entezari_arxiv_2022 conjectured that, once permutation symmetries of hidden units are taken into account, LMC should hold with high probability. Building on this, Ainsworth_ICLR_2023 introduced a weight-matching approach based on bipartite graph alignment, while Pena_CVPR_2023 applied Sinkhorn’s algorithm as a relaxation to solve the alignment problem more directly. Although a number of papers (Luca_JMLR_2019; Nguyen_ICLR_2019; Nguyen_ICML_2019; Kuditipudi_NIPS_2019) have studied nonlinear mode connectivity, theoretical understanding of LMC remains limited. Ferbach_AISTATS_2023 established width-dependent conditions under which LMC can be guaranteed, assuming independence of weight vectors. Zhou_NIPS_2023 proposed the notion of layerwise linear feature connectivity (LLFC) and proved that LLFC implies LMC. More recently, Ito_ICLR_2025 highlighted the role of dominant singular vectors in parameter space, showing that they are critical to satisfying LMC, especially when alignment via weight matching is applied.
In addition, several theoretical works have analyzed the effect of network width on nonlinear connectivity. For instance, Nguyen_ICML_2019 proved that in over-parameterized neural networks with piecewise-linear activations, every sublevel set of the loss is connected and unbounded. Shevchenko_ICML_2020 further showed that in wide multilayer perceptrons, SGD solutions can be linked through piecewise-linear paths along which the loss barrier vanishes as width increases. However, to the best of our knowledge, there are no theoretical results directly establishing that larger width makes LMC more likely. This motivates our empirical investigation of that question.
@@
Model merging has been studied in close connection with LMC, as well as in the context of federated learning. The idea of federated learning was introduced by McMahan_AISTATS_2017 and Konecny_arxiv_2016, where models are trained locally on partitioned datasets. Wang_ICLR_2020 explored permuting local model components prior to aggregation, while Singh_NIPS_2020 proposed an approach to merge models by optimal transport, conceptually related to the weight-matching method of Ainsworth_ICLR_2023. Although the method of Signh_HiLD_2024 was primarily designed for fusion and empirically underperforms weight matching, it can still be interpreted as an LMC-based technique because it enforces alignment within the same architecture.
While our work focuses on merging models trained from different random seeds, there is a parallel line of research on merging models obtained by fine-tuning from the same fundamental model. A representative example is Model Soups (Wortsman_ICML_2022), which showed that averaging weights across fine-tuned models trained under different hyperparameters can improve test accuracy without additional inference cost, contrasting with standard ensembling. Matena_NeurIPS_2022 extended this idea by weighting models according to the Fisher information matrix, leading to more effective combinations. Yadav_NeurIPS_2023 introduced TIES-Merging, which prunes parameters with small updates, resolves conflicting signs, and averages only aligned components. These approaches are primarily designed for merging fine-tuned models of the same pretrained model, not for models independently trained from scratch with different initializations. By contrast, our study addresses this more challenging setting and seeks to shed light on the feasibility of model merging under such conditions.
Relation between Model Width and Loss Landscape.
A broad trend in prior work is that increasing width tends to simplify the loss landscape of deep neural networks and makes the set of optimal solutions more connected. Nguyen_arxiv_2021; Simsek_ICML_2021 pointed out that the topology of the optimal solution set can change drastically by adding only a small number of neurons. Specifically, Nguyen_arxiv_2021 proved that for deep networks, having a single wide layer of width $n+1$, where $n$ is the number of training samples, is sufficient to ensure that all sublevel sets of the training loss are connected. Simsek_ICML_2021 further showed that adding one extra neuron per layer is enough to turn discrete minima into a single connected manifold. Liang_NeurIPS_2018 showed that adding one special neuron to each layer can eliminate all bad local minima, so that every local minimum becomes a global minimum. Subsequent works have also reported phase-transition-like behavior as width increases. For example, Li_SIAM_2022 showed that there exists a critical width $m^{\ast}$ such that when $m\geq m^{\ast}$, all suboptimal basins disappear and only optimal basins remain. More recently, Kim_ICLR_2025 established that in regularized two-layer networks, once the width exceeds a certain threshold, the optimal solution set becomes connected in parameter space. These results collectively suggest that widening changes the structure of the optimal set and can eventually yield mode connectivity, especially under regularization. Building on this line of work, we empirically show that, after calibrating the loss via inverse temperature scaling, the stronger property of linear mode connectivity holds in sufficiently wide networks without permutations. We believe our findings point to new directions for understanding SGD learning dynamics and may also inform future techniques for model merging and federated learning.
@@ Appendix B Experimental Settings
This section details the setup used to train neural networks and obtain solutions from stochastic gradient descent (SGD). The experiments were conducted on four benchmark datasets: MNIST (MNIST), Fashion-MNIST (FMNIST) (FMNIST), CIFAR-10 (CIFAR10), and CIFAR-100. For weight matching (WM), we adopted the implementation available in the public repository of Ainsworth_ICLR_2023333https://github.com/samuela/git-re-basin.
All training and evaluation procedures were performed on a Linux workstation equipped with two AMD EPYC 7543 32-core processors, eight NVIDIA A30 GPUs, and 512 GB of memory. We used PyTorch 2.5.1444https://pytorch.org/, PyTorch Lightning 2.4.0555https://lightning.ai/docs/pytorch/stable/, and torchvision 0.20.1666https://pytorch.org/vision/stable/index.html as the software framework.
@@ B.1 Model Training
In line with the setup of Ainsworth_ICLR_2023, we trained a fully connected multilayer perceptron (MLP) with three hidden layers, each containing 512 units. ReLU activations were applied to the hidden layers. For both MNIST and FMNIST, training used the Adam optimizer with a fixed learning rate of $1\times 10^{-3}$ and a weight decay of $3\times 10^{-3}$. The batch size was 512, and models were trained for up to 100 epochs. No learning rate scheduling was applied.
@@
For these experiments, we employed the source code released by Ito_ICML_2025777https://github.com/e5-a/STE-MM. The VGG-11 and ResNet-20 architectures followed the implementations described by Ainsworth_ICLR_2023. During model merging, BatchNorm statistics were re-calibrated using the training data, following the procedure of Jordan_ICLR_2023. Models were optimized with SGD using a learning rate of $0.4$ and a weight decay of $3\times 10^{-3}$. The batch size was set to 500, and training was carried out for a maximum of 100 epochs. Data augmentation included random $32\times 32$ crops and random horizontal flips.
@@ Appendix C Layerwise Linear Feature Connectivity
Zhou_NIPS_2023 pointed out that the existence of LMC in settings such as permutation and spawning can be attributed to a property called layerwise linear feature connectivity (LLFC). Additionally, Zhou_NIPS_2023 also showed that two conditions are sufficient for LLFC to hold: weak additivity for ReLU activations (Definition 5.1) and the commutativity property defined as follows:
@@
C.1 Difference between LLFC and LEWC
We empirically confirmed that, in LMC without permutations, LLFC does not hold whereas LEWC does. In this subsection, we provide a qualitative explanation for why LLFC fails even though LEWC holds.
A key difference between LLFC and LEWC is that LLFC is a property that arises when the weights of two models are sufficiently close, whereas LEWC is a property that arises when the weights of two models are different, namely approximately orthogonal. This can be understood by focusing on the difference between commutativity and reciprocal orthogonality, which are the respective sufficient conditions that distinguish LLFC from LEWC.
Below we explain this point more concretely using equations. Let be $\bm{\theta}_{a}$ and $\bm{\theta}_{b}$ the two models, and for simplicity assume that biases can be ignored. Commutativity for $\ell\geq 1$ requires
$\bm{W}_{\ell}^{(a)}\bm{z}_{\ell-1}^{(a)}+\bm{W}_{\ell}^{(b)}\bm{z}_{\ell-1}^{(b)}=\bm{W}_{\ell}^{(a)}\bm{z}_{\ell-1}^{(b)}+\bm{W}_{\ell}^{(b)}\bm{z}_{\ell-1}^{(a)}\Leftrightarrow(\bm{W}_{\ell}^{(a)}-\bm{W}_{\ell}^{(b)})(\bm{z}_{\ell-1}^{(a)}-\bm{z}_{\ell-1}^{(b)})=0.$
This condition clearly holds when all weights match, that is, when $\|\bm{W}_{\ell}^{(a)}-\bm{W}_{\ell}^{(b)}\|=0$. Zhou_NIPS_2023 stated that this explains why WM can find permutations that make LMC hold, since the objective of WM is to search for permutations that minimize the distance between the weights of two models. In spawning, it is also known that LMC holds between solutions obtained by branching from a partially trained initialization and then continuing SGD independently. This is because the first training steps before branching keep the two models close, which promotes commutativity. Zhou_NIPS_2023 experimentally verified that commutativity holds in the spawning setting as well. In contrast, reciprocal orthogonality clearly does not hold when the two weight matrices match. Specifically, it requires $\bm{W}_{\ell}^{(a)}\bm{z}_{\ell}^{(b)}=0$ and $\bm{W}_{\ell}^{(b)}\bm{z}_{\ell}^{(a)}=0$, but if the weight matrices are equal and biases are negligible, then the activations are also equal, and these conditions are not satisfied. This shows that commutativity requires the two models to be sufficiently close, whereas reciprocal orthogonality requires them to be sufficiently different, namely approximately orthogonal, so the two requirements are incompatible. For example, in Figures 11(a) and 7 we showed that without permutations, commutativity fails while reciprocal orthogonality holds, which reflects that the two models remain sufficiently distinct. Conversely, when models are aligned using WM discovered permutations, or in the spawning setting, the weight matrices are expected to be close. In such cases, commutativity is likely to hold, whereas reciprocal orthogonality is unlikely to be satisfied.
This incompatibility implies that the reason for LMC identified in our paper is fundamentally different from the LLFC based explanation of LMC in Zhou_NIPS_2023. Intuitively, under commutativity, permutations make the two models sufficiently close, so the merged model also remains close to each of the original models, and its output stays similar to theirs. In contrast, in our explanation, the two models have weight matrices that are highly different, namely approximately orthogonal, which allows the merged model to embed the two functions independently.
This distinction also explains why inverse temperature calibration is necessary under LEWC. LLFC requires the two weight matrices to be close, so the norm of the merged weight matrix is similar to that of the original models. In contrast, LEWC requires the two weight matrices to be orthogonal, so the norm of the merged weight matrix is strictly smaller than that of each original model. Since this reduction occurs at every layer, the output norm decays exponentially with depth. This provides the reason for the exponential decay of activation norms under LEWC.
Appendix D Relation to Loss Landscape Flatness
(a) Test accuracy of the perturbed model $\bm{\theta}_{a}+\lambda d\bm{u}/\|\bm{u}\|$ obtained by moving $\bm{\theta}_{a}$ along a random direction.
(b) Uncalibrated test loss of the perturbed model $\bm{\theta}_{a}+\lambda d\bm{u}/\|\bm{u}\|$ obtained by moving $\bm{\theta}_{a}$ along a random direction.
Figure 12: Test accuracy and test loss of models perturbed along a random direction.
We showed that LMC becomes easier to satisfy as the width increases. One possible alternative explanation is that widening simply improves the flatness of the loss or accuracy landscape, and that LMC emerges as a consequence of this increased flatness. In this section, we examine this possibility and show that a mere improvement in flatness is insufficient to explain our observations.
D.1 Empirical Evaluation of Flatness
As a simple proxy for flatness with respect to width, we evaluate how test accuracy and test loss change when we perturb a trained model along a random direction. Let $\bm{\theta}_{a}$ and $\bm{\theta}_{b}$ be two trained models. We first compute their $L^{2}$ distance $d=\|\bm{\theta}_{a}-\bm{\theta}_{b}\|$. Although $d$ could in principle grow with width, in our setting weight decay prevents it from increasing substantially. We then generate a Gaussian random vector $\bm{u}$ and normalize it to unit length, and use a perturbation of the same norm as the inter model distance, namely $d\bm{u}/\|\bm{u}\|$. This construction is intended to compare random perturbations at the same scale as the distance between two independently trained models. For each $\lambda\in[0,1]$, we measure the accuracy and loss of the perturbed model $\bm{\theta}_{a}+\lambda d\bm{u}/\|\bm{u}\|$. We expect performance to degrade as $\lambda$ approaches 1, and we interpret the rate of degradation as a proxy for flatness. The results are shown in Figure 12. As the width increases, both test accuracy and test loss become almost insensitive to the perturbation. In particular, even at $\lambda=1$, the degradation becomes negligible after a modest increase in width. This indicates that this flatness proxy improves rapidly with width.
Figure 13: Cosine similarity of hidden layer activations between $\bm{\theta}_{a}$ and the perturbed model $\bm{\theta}_{a}+d\bm{u}/\|\bm{u}\|$.
D.2 Why Does the Flatness Improve with Width
Intuitively, one might expect that moving model parameters in a random direction would quickly harm accuracy and loss. However, Figure 12 shows that widening significantly suppresses this degradation, even more so than in the interpolation between two independently trained models shown in Figures 1 and 2. We attribute this to the nature of the random perturbation. When the width is large, a perturbation with fixed overall norm is spread over a much larger number of parameters, so the per parameter magnitude becomes relatively small. As a result, the perturbed model behaves similarly to the original one.
More formally, let $\bm{W}_{\ell}$ denote the weight matrix of the original model at layer $\ell$, and let $\bm{W}^{\prime}_{\ell}$ denote the random perturbation added to that layer. The perturbed weights are $\bm{W}_{\ell}+\bm{W}^{\prime}_{\ell}$. Ignoring biases, the pre-activation becomes $(\bm{W}_{\ell}+\bm{W}^{\prime}_{\ell})\bm{z}_{\ell-1}=\bm{W}_{\ell}\bm{z}_{\ell-1}+\bm{W}^{\prime}_{\ell}\bm{z}_{\ell-1}$. Thus, the change in output induced by the perturbation is represented by $\bm{W}^{\prime}_{\ell}\bm{z}_{\ell-1}$. The $i$ th component of this term is $\sum_{j}\bm{W}^{\prime}_{\ell,i,j}\bm{z}_{\ell-1,j}$. Since $\bm{W}^{\prime}_{\ell}$ is independent of $\bm{z}_{\ell-1}$ and its total norm is fixed to $d$, the typical magnitude of each entry $\bm{W}^{\prime}_{\ell,i,j}$ decreases as the width increases, because the same norm is distributed across more parameters. Moreover, due to the effect of weight decay, the inter model distance $d$ does not necessarily grow with width. Therefore, $\sum_{j}\bm{W}^{\prime}_{\ell,i,j}\bm{z}_{\ell-1,j}$ is expected to approach zero as the width increases.
To validate this intuition, we measure the cosine similarity of hidden layer activations between the original model $\bm{\theta}_{a}$ and the perturbed model $\bm{\theta}_{a}+d\bm{u}/\|\bm{u}\|$. The results are shown in Figure 13. The similarity increases with width, supporting the view that the apparent improvement in flatness under random perturbations is largely explained by the diminishing effect of Gaussian noise in wide networks.
Figure 14: Cosine similarity of hidden layer activations between $\bm{\theta}_{a}$ and $\bm{\theta}_{b}$.
D.3 Flatness alone Cannot Explain Our Findings
As a comparison to the random perturbation setting, we examine the cosine similarity of hidden layer activations between two independently trained models $\bm{\theta}_{a}$ and $\bm{\theta}_{b}$, shown in Figure 14. Even though the norm of the random perturbation in Figure 13 equals the model distance between $\bm{\theta}_{a}$ and $\bm{\theta}_{b}$, the activation similarities exhibit qualitatively different behavior. This suggests that probing flatness via random perturbations does not capture the relationship between independently trained models produced by SGD.
Finally, note that if LMC were explained solely by increased flatness, then the exponential decay of activation norms predicted by LEWC would also need to follow from flatness, which it does not. Moreover, if flatness were the only reason, the merged model should achieve test loss comparable to the original models even without inverse temperature calibration. However, Figure 2 contradicts this. These observations indicate that improved flatness alone is insufficient to account for the emergence of LMC with increasing width.
Appendix E Proof
E.1 Main Theorem
Theorem E.1.